OpenStudy (anonymous):

Boolean Algebra: Reduce literals (Eqations) Hello, can someone please provide me with how to solve this problem? I must reduce this expression: a'b'c' + a'bc' + a'bc + ab'c' the solution should be b'c' + a'b Please explain the axioms used.

OpenStudy (anonymous):

This site is seriously lacking community help in Physics and Computer Science. This would be my second post where I have answered my own question. a'b'c' + a'bc' + a'bc + ab'c' Use Distributive to get: a'c'( b' + b ) + a'bc + ab'c' Identity : ( b' + b ) = 1 Group like terms and use Distributive Prop. a'( c' + cb ) + ab'c' Use ID (a + a'b = a + b) where a=c' and b=b to get: a'( c' + b ) + ab'c' Redistribute the terms and group c' and repeat the process to get: c'a' + c'b' + a'b I hope we get more community support in these tougher topics.

OpenStudy (anonymous):

i agree